OpenStudy (pokemon23):

9+2(5v+3)=13v show all work and check Saifoo does it again :) alright

OpenStudy (anonymous):

9+10v+6=13v

OpenStudy (anonymous):

can you solve it from there?

OpenStudy (anonymous):

i think it should be +6 not +3 saifoo?

OpenStudy (pokemon23):

too lazy doggie I'm pissed at my teacher earth science

OpenStudy (saifoo.khan):

I know. i did wrong. i was checking u. LOL.

OpenStudy (anonymous):

haha sorry what she/he do

OpenStudy (pokemon23):

lol saifoo

OpenStudy (saifoo.khan):

9 + 10v + 6 = 13v 10v - 13v = -6-9 -3v = -15 v = 15/3 v = 5